a batter hits a pop-up that remains in the air for 7s before hitting the ground. how fast was the ball travelling when it returned to the ground?
1
Expert's answer
2020-11-10T06:57:47-0500

"v=g\cdot\frac t2=9.8\cdot \frac 72 =34.3" ("\frac ms" ).